OpenStudy (anonymous):

Find the four arithmetic means between -21 and -36.

OpenStudy (anonymous):

maybe this example can help: http://www.algebra.com/algebra/homework/equations/Equations.faq.question.332361.html Sorry, if it doesn't.

OpenStudy (anonymous):

make it to (n+1) divisions.. \[d = \frac{b-a}{n+1}=\frac{-36-(-21)}{4+1}\] and keep adding this d to -21 to get your numbers. adding the fifth time should give you -36

OpenStudy (anonymous):

-24,-27,-30 and -33